### v1

# Non-Relativistic Hyperboloidal Method for Quasinormal Modes

Abstract
The recently reported compactified hyperboloidal method has found wide use in the numerical computation of quasinormal modes, with implications for fields as diverse as gravitational physics and optics. We extend this intrinsically relativistic method into the non-relativistic domain, demonstrating its use to calculate the quasinormal modes of the Schrödinger equation and solve related bound-state problems. We also describe how to further generalize this method, offering a perspective on the importance of non-relativistic quasinormal modes for the programme of black hole spectroscopy.
I Introduction
Quasinormal modes (QNMs) are complex frequency modes which characterize the resonant response of a system to linear perturbations. They are prevalent in the physics of waves, with special prominence in optics and gravitational physics. In optics, QNMs are useful for understanding the behaviour of resonant photonic structures, such as plasmonic crystals, nanoparticle traps, metal gratings, and optical sensors Lalanne2018 ; Yan2018 ; Qi2021 ; Gras2019 ; Ren2021 . In gravitational physics, they are thought relevant to tests of black hole no-hair conjectures Gossan2012 ; Shi2019 ; Ma2023 , and central to the emerging project of black hole spectroscopy with gravitational waves Dreyer2004 ; Cabero2020 . While the QNM literature in optics treats dispersion as a matter of necessity lalanne2019 ; Primo2020 , the prevailing methods in gravitational physics are concerned with non-dispersive, relativistic wave propagation Kokkotas1999 ; Berti2009 ; Konoplya2011 . We believe there are good reasons to go beyond relativistic wave propagation in the gravitational context. A variety of quantum gravity models predict the dispersive propagation of gravitational waves Nishizawa2018 ; Mastrogiovanni2020 ; Ezquiagal2021 ; Aydogdu2022 , for example, in models with a non-zero graviton mass, violation of Lorentz invariance, and higher dimensions Will1998 ; Sefiedgar2011 ; Kostelecky2016 . Indeed, it has been proposed that QNMs may be used to probe gravity beyond general relativity, through imprints on radiative emission from black holes Glampedakis2019 ; Agullo2021 ; Srivastava2021 ; Chen2021 ; Fu2024 . More generally, we anticipate that development of QNM methods for non-relativistic operators will broaden the scope of existing questions in QNM theory.
Numerical methods underpin much of the progress in QNMs over recent years. Indeed, efficient schemes for computing the QNMs of potentials are likely indispensable for future developments in both theory and the modelling of observations. Recently, the so-called compactified hyperboloidal method has proven to be a powerful tool Jaramillo2021 , finding wide use in the computation of black hole QNM spectra and bringing within reach the systematic exploration of their connection to pseudospectra Destounis2021 ; Ripley2022 ; Gasperin2022 ; Sarkar2023 ; Boyanov2024 ; Cao2024 ; Zhu2024 ; Destounis2024 . Beyond this, it is natural to ask whether the method can also find use in optical systems. We believe it can, but it cannot be widely applied in optics without modification. This is because optical media create non-relativistic and dispersive dynamics, while the present formulation of the method treats only relativistic and non-dispersive dynamics, as may be seen from its use of hyperbolic spatial slices penetrating the black hole horizon and future null infinity.
A notable optical system that motivates the development of a hyperboloidal method for optics is the fiber optical soliton, which has recently been established as a black hole analogue with an exactly known QNM spectrum Burgess2024 . As such, the soliton is the ideal system with which to develop the method, as the resulting numerics can be compared both to known analytical results and to the numerics of the corresponding relativistic system. Moreover, perturbations to the soliton realize the Schrödinger equation with a Pöschl-Teller potential, making the soliton a promising experimental platform with which to address questions in QNM theory, such as the physical status of spectral instabilities observed in QNM numerics, where the Pöschl-Teller potential is paradigmatic Ferrari1984 ; Nollert1993 ; Cho2012 ; Jaramillo2021 ; Cardoso2024 .
In this article, we outline a new method for the numerical computation of QNM spectra for operators with a non-relativistic dispersion relation, by adapting the compactified hyperboloidal method. We begin by showing how to compute the QNMs of the Schrödinger equation for an arbitrary potential, noting that the relativistic and non-relativistic spectra are related by a simple endomorphism. We subsequently demonstrate the method for the Pöschl-Teller potential, explicitly calculating the soliton QNM spectrum numerically. Finally, we sketch how to develop these ideas in order to treat generalized non-relativistic dispersion relations, and discuss potential applications of the more general method, with emphasis on its future use in black hole spectroscopy.
II Compactified Hyperboloidal Method for the Schrödinger Equation
We begin by considering a scalar field $\phi$ which obeys a Schrödinger equation of the form,
$(i\partial_{t}-\partial_{r}^{2}+V)\phi=0,$ (1)
with $V$ a potential that vanishes for $r\to\pm\infty$. In order to compute the QNMs of this equation, we first transform from coordinates for which the QNMs diverge at the spatial boundaries, to ones for which they remain finite. Following Jaramillo2021 , we adopt coordinates,
$t=\tau-h(x),\quad r=g(x),$ (2)
where $g(x),h(x)$ are yet to be given, and $\partial_{\tau}=\partial_{t}$ by construction. In the relativistic method, $h(x)$ is chosen so that contours of $\tau$ tend asymptotically to the horizon and light cone. However, there is no preferred speed in our system, and we parameterize $h(x)$ to obtain hyperbolic slices propagating outwards with asymptotic speed $v_{a}$ in the $(r,t)$ coordinates. In particular, we write
$g(x)=\arctan{x},\quad h(x)=v_{a}^{-1}\log(1-x^{2}),$ (3)
where $g(x)$ compactifies the space such that the real line of $r$ gives $x\in[-1,1]$ if we close the set by including the boundaries. By considering mode solutions far from the potential, it is easily confirmed that those QNMs whose contours of constant amplitude propagate outwards with speed $v_{a}$ take finite amplitudes on the new spatial boundaries. However, dispersion in this non-relativistic system implies that asymptotic contours of constant amplitude do not generally align with those of constant phase, meaning finite amplitudes on the boundaries are typically accompanied by phase singularities. This is resolved by a phase-rotation of the field,
$\hat{\phi}=e^{-\Delta\log(1-x^{2})/2}\phi,$ (4)
with $\Delta=i(v_{a}-v_{p})/2$ and $v_{p}$ the asymptotic phase velocity associated with $\phi$. The form of the phase rotation follows from the asymptotic dispersion relation of Eq. (1), and intuitively depends on the mismatch of the two speeds. This ensures $\hat{\phi}$ is well-defined on the compactified space, but its construction presumes two real parameters, $v_{a}$ and $v_{p}$. Indeed, for mode solutions, these parameters are characteristic, in that they determine the frequency of the mode. In particular, we find
$\Omega=-\frac{1}{2}v_{a}\bigg{(}v_{p}+i\sqrt{v_{a}(v_{a}-2v_{p})}\bigg{)},$ (5)
having used $\Omega=-K^{2}$, the asymptotic dispersion relation of Eq. (1), where $K$ is the asymptotic wavenumber in the $(t,r)$ coordinates. That we can express the mode frequency in terms of asymptotic speeds is a direct consequence of dispersion. In a relativistic system, all asymptotic speeds are the speed of light, and $\Omega$ cannot be expressed in terms of asymptotic speeds. This difference between the relativistic and non-relativistic methods is crucial, as will be explained in what follows.
We proceed as in Jaramillo2021 , by rewriting Eq. (1) in the new coordinates and performing a first-order reduction in time, introducing the auxiliary field $\hat{\psi}$. The equation of motion becomes $\partial_{\tau}\hat{\phi}=\hat{\psi}$ with
$x^{2}\partial_{\tau}\hat{\psi}=J_{1}\hat{\phi}+J_{2}\hat{\psi},$ (6)
where $J_{1}$ and $J_{2}$, given in the Appendix, are spatial operators depending on the potential and the asymptotic speeds. In contrast to the relativistic method, $\partial_{\tau}\hat{\psi}$ cannot be isolated by division in Eq. (6) because its pre-factor vanishes at $x=0$. This unavoidably occurs as a consequence of the hyperboloidal slices necessarily “turning around” in order to describe contours that are outgoing in both directions. Our alternative approach is to construct the $\partial_{\tau}\hat{\psi}$ using a Taylor series of its spatial derivatives at $x=0$, which may be obtained by repeated differentiation of Eq. (6). In fact, this treatment is necessary only for those terms indivisible by $x^{2}$, and we obtain a simpler result if we initially separate the terms in this way. This separation is mostly trivial, but for the potential we write $V=V_{0}+xV_{1}+x^{2}\tilde{V}(x)$, where $V_{0}$ and $V_{1}$ are Taylor series coefficients about $x=0$ and $\tilde{V}$ accounts for the remaining terms. We obtain
$\partial_{\tau}\hat{\psi}=L_{1}\hat{\phi}+L_{2}\hat{\psi},$ (7)
where $L_{1}$ and $L_{2}$ are spatial operators that we derive in the Appendix. Eq.(7) is formally identical to that obtained in the relativistic method Jaramillo2021 , but the operators are quite different, containing arbitrarily high spatial derivatives and depending on the asymptotic speeds.
In matrix form, we write $i\partial_{\tau}u=Lu$ with
$u\equiv\begin{pmatrix}\hat{\phi}\\ \hat{\psi}\end{pmatrix},\quad L\equiv i\begin{pmatrix}0&1\\ L_{1}&L_{2}\end{pmatrix},$
and obtain the mode equation
$Lu=\Omega u.$ (8)
Importantly, the operator $L$ in the above equation is actually parameterized by the mode frequency $\Omega$. This is because the operator depends on $v_{a}$ and $v_{g}$, which may be expressed in terms of $\Omega$ and its complex conjugate using Eq.(5). This is an essential difference from the relativistic method, where the corresponding operator contains no dependence on $\Omega$ Jaramillo2021 . Nevertheless, Eq.(8) unambiguously determines the QNM spectrum.
Eq (8) is discretized using $N$-point Chebyshev nodes of the second kind. In this way, fields are approximated by $N$-dimensional vectors and spatial operators by $N$-dimensional matrices. It follows that the vector $u$ and the operator $L$ are approximated by $2N$-dimensional vectors and matrices, respectively. The result is
$L^{N}u^{N}=\Omega\ \!u^{N}.$ (9)
The QNM spectrum may then be obtained in the usual way using $\det(L^{N}-\Omega\ \!\text{Id})=0$. In the Appendix, we show that this determinant may be rewritten as that of a smaller $N$-dimensional matrix, $M$. Its elements are quadratic in the square root of the QNM frequency, giving rise to a polynomial of degree $2N$ in $\sqrt{\Omega}$. For a given potential $V$, the roots may be numerically determined in order to give $2N$ of the QNM frequencies. The fact that the frequency enters via its square root is a result of the Schrödinger equation having a first derivative in time, rather than a second derivative in time like the wave equation. Indeed, the exact QNM spectra of the Schrödinger and wave equations are related to each other by $i\sqrt{\Omega}=\omega$, as elaborated in Burgess2024 . This means we can relate the results of the non-relativistic method to those of the relativistic method, allowing us to better evaluate the accuracy of the new method.
III Quasinormal Modes of the Pöschl-Teller Potential
In this section, we use the above numerical method to calculate the QNMs of the Schrödinger equation with the Pöschl-Teller potential,
$V=V_{0}\ \!\text{sech}^{2}(r)=V_{0}(1-x^{2}),$ (10)
which serves as an exemplar for both the relativistic and non-relativistic methods. Its QNMs are finite polynomials in the compactified spatial coordinate, with the result that an $N$-point discretization reproduces the first $2N$ QNMs to arbitrary precision. The Pöschl-Teller potential is also ideal because the corresponding QNM spectrum of the Schrödinger equation is given analytically by
$\Omega_{n}=\Bigg{[}\!\ n+\frac{1}{2}-i\sqrt{V_{0}-\frac{1}{4}}\!\ \Bigg{]}^{\!2}\!\!,$ (11)
allowing us to verify our results Burgess2024 . In regards to the non-relativistic method, we note that the Pöschl-Teller potential is especially simple because all its QNMs have the same $v_{a}$ parameter, which is a result of the unique fact that $i\sqrt{\Omega}$ is aligned along vertical lines in the complex plane for this potential. Lastly, we partition the Pöschl-Teller potential with $V_{1}=0$ and $\tilde{V}=-V_{0}$, which reinforces the simplicity of the potential.
Now, we make some comments on the specifics of our implementation of the method. We find the calculation is significantly more efficient for odd $N$. This is a consequence of discretization. The Taylor series expansions of $L_{1}$ and $L_{2}$ involve spatial derivatives at $x=0$, which are obtained by integration with a Dirac delta function in the continuous case, and by matrix multiplications in the discretized case. For odd $N$, the relevant matrix is zero everywhere but a central column whose entries are unity. However, for even $N$, the matrix is everywhere populated, and this increases the computational cost of the calculation. We also find that evaluating the determinant of the large symbolic matrix $M$ is inefficient, so we instead sample the determinant in the complex plane and reconstruct the symbolic determinant using polynomial interpolation. This uses that the method produces a polynomial of degree $2N$ in $\sqrt{\Omega}$. Importantly, this is true no matter what potential we consider.
Figure 1: QNM spectra for the Schrödinger equation with a potential $V=V_{0}\hskip 1.5pt\text{sech}^{2}(r)+\epsilon\Delta V$, where $V_{0}=1$ and $\Delta V$ is the perturbation shown in green in the inset, alongside the unperturbed Pöschl-Teller potential in blue. The red crosses and blue circles correspond to the unperturbed Pöschl-Teller potential ($\epsilon=0$), with red crosses given by the exact formula and blue circles numerically determined by the new method. The green diamonds correspond to a perturbed potential ($\epsilon=10^{-30}$) and are also numerically determined. (a) displays the three QNM spectra, while (b) displays the same spectra under the transformation $\Omega\to i\sqrt{\Omega}$, which relates the spectra to those of a corresponding relativistic operator.
In Figure 1(a), we plot the exact QNM frequencies of the unperturbed Pöschl-Teller potential Burgess2024 alongside those calculated by the new numerical method, with a resolution of $N=201$. We find excellent agreement for all frequencies, with an error which may be made arbitrarily small by increasing the working precision. These results are given in Figure 2. We also calculate the QNM spectrum of a perturbed Pöschl-Teller potential,
$V=V_{0}(1-x^{2})+\epsilon\Delta V,$ (12)
where $\epsilon=10^{-30}$ and $\Delta V$ is a randomly chosen polynomial of degree 9, shown in the inset of Figure 1. We find that the spectrum closely resembles the unperturbed spectrum up to the $10^{\text{th}}$ overtone index, beyond which the frequencies are significantly displaced from their unperturbed values, as shown in Figure 2. These numerical results are then indicative of spectral instabilities that have been reported by previous authors Nollert1996 ; Daghigh2020 ; Jaramillo2021 .
Figure 2: Comparisons of exact and numerically determined QNM frequencies for the Schrödinger equation with a potential $V=V_{0}\hskip 1.5pt\text{sech}^{2}(r)+\epsilon\Delta V$, where $V_{0}=1$ and $\Delta V$ is the perturbation shown in Figure 1. The first 20 QNM frequencies are displayed. The unperturbed ($\epsilon=0$) spectrum is recovered well by the new numerical method, with errors smaller than $10^{-18}$ for the chosen working precision. We also obtain the perturbed ($\epsilon=10^{-30}$) spectrum and find the deviation from the exact spectrum grows radidly with overtone index, $n$, as in previous works on spectral instability.
The simple relationship between the QNMs of the Schrödinger and wave equations becomes visible under the transformation $\Omega\to i\sqrt{\Omega}$, which maps the spectrum of the former onto that of the latter. In Figure 1(b), we plot $i\sqrt{\Omega}$ for the same spectra as above, obtaining the recognizable vertical lines in the complex plane that are characteristic of the wave equation with a Pöschl-Teller potential. In this way, we illustrate how one can cross-verify the results of the relativistic and non-relativistic methods against each other, for arbitrary potentials.
IV Discussion
In this section, we discuss potential applications of the non-relativistic compactified hyperboloidal method that we developed in the preceding text, suggesting well-motivated directions in which to further develop the method and providing a sketch of how this can be achieved. The principal application of the method is the calculation of QNM frequencies of the Schrödinger equation equipped with a general potential. Using our method, we numerically calculated QNM spectra for the Pöschl-Teller potential and perturbations of that potential, finding agreement with earlier works Boonserm2011 ; Cardona2017 ; Burgess2024 . In addition, we note that this method may also be used to numerically solve for the quantum mechanical bound-states of a general potential well, using the well-known connection between the QNMs of a potential barrier and the bound-states of the corresponding well Ferrari1984b .
The non-relativistic method we have presented readily generalizes beyond the Schrödinger equation, allowing us to treat a large class of more general non-relativistic operators. Indeed, the method presented in this paper primarily serves a didactic purpose, as a demonstration of a general approach with which one may calculate QNMs of these more general operators. The primary motivation for this is to facilitate the efficient computation of QNMs of operators that deviate from the wave equation only by the presence of weak dispersion, as are known to arise in models of quantum gravity, where a thoroughgoing understanding of QNMs is of special interest. The modelling of dispersive gravitational wave propagation and its influence on the observable QNM spectrum will be essential if black hole spectroscopy is to be an effective probe into the domain of quantum gravity.
A further motivation for generalizing the non-relativistic method is to shed light on QNM spectral instabilities, and facilitate experimental tests of the recent ultraviolet universality conjecture, which posits that sufficiently high overtones converge to logarithmic Regge branches in the complex plane, in the high-frequency limit of potential perturbations Jaramillo2021 ; Boyanov2024 . This effect is easily seen in numerical calculations of the Pöschl-Teller spectrum, on account of its simplicity, but has yet to be experimentally confirmed. Using the optical soliton, whose perturbations realize this potential, an experimental test becomes possible. However, interpreting observations of the migration of soliton QNM frequencies demands an understanding of the higher-order dispersive effects which will be present in any experiment. In particular, the influence of weak third-order dispersion acting on the perturbative probe field should be incorporated into the analysis, in order to provide the best test of the above conjecture. This motivates the development of the non-relativistic method beyond the Schrödinger equation, to include higher-order dispersive terms.
In view of the above reasons to generalize the non-relativistic method, we present a sketch of the more general method, which we will elaborate in future work. Suppose we have a non-relativistic equation of the form
$(\alpha(i\partial_{t})+\beta(i\partial_{r})+V)\phi=0,$ (13)
with $\alpha(z)$ and $\beta(z)$ finite polynomials in $z$, and $d$ the larger degree among the two polynomials. We apply a hyperboloidal coordinate transformation and a phase rotation of the fields, parameterised by the asymptotic speeds $v_{a}$ and $v_{p}$. Then, we introduce auxiliary fields to effect a $d$th-order reduction in time, defining
$\phi_{1}=\phi,\quad\phi_{k+1}=\partial_{\tau}\phi_{k},$ (14)
with $1\leq k<d$. This closely mirrors the treatment of resonator QNMs in optics lalanne2019 . The equations of motion of these fields are trivial for all fields but $\phi_{d}$, whose equation of motion more closely resembles Eq. (6). If we use a Taylor series expansion of $\partial_{\tau}\phi_{d}$ around zero, we can write it in terms of spatial operators acting on the fields. The general form of the now $d$-dimensional operator $L$ is
$L=\begin{pmatrix}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ 0&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&1\\ L_{1}&L_{2}&L_{3}&\cdots&L_{d}\\ \end{pmatrix},$ (15)
which we discretize as before. Then, we use the asymptotic dispersion relation of Eq. (13) to eliminate the asymptotic speeds, obtaining a vector equation for the QNM frequencies. From the form of $L$, it can be shown that it is always possible to construct an $N$-dimensional matrix $M$ whose determinant is a finite polynomial for the QNM frequencies. This may then be solved numerically and the frequencies $\Omega$ determined.
Future works can also develop this non-relativistic method, along the lines of Jaramillo2021 , in order to calculate the pseudospectra of non-relativistic operators. It is our view that the relationship between perturbed QNM spectra and the pseudospectrum is best understood from a broader perspective, not limited to relativistic wave operators. Lastly, we expect that numerical methods will become increasingly important for addressing questions in QNM theory, and we anticipate that investigations into the QNMs of non-relativistic fields will provide new avenues to explore these questions.
The supporting data for this Letter are openly available from data_source .
We would like to express our thanks to Théo Torres for providing us a useful overview at the outset of this research. This work was supported in part by the Science and Technology Facilities Council through the UKRI Quantum Technologies for Fundamental Physics Programme [Grant ST/T005866/1]. CB was supported by the UK Engineering and Physical Sciences Research Council [Grant EP/T518062/1].
For the purpose of open access, the authors have applied a creative commons attribution (CC BY) licence (where permitted by UKRI, ‘open government licence’ or ‘creative commons attribution no-derivatives (CC BY-ND) licence’ may be stated instead) to any author accepted manuscript version arising.
Appendix A Appendix
A.1 1. Equation of Motion for the $\hat{\psi}$ Field
The auxilliary field $\hat{\psi}$ obeys Eq. (6), with
$J_{1}$ $=v_{a}^{2}\Big{[}V+\Delta(1-x^{2})-\Delta^{2}x^{2}+2(1+\Delta)x(1-x^{2})\partial_{x}-(1-x^{2})^{2}\partial_{x}^{2}\Big{]},$
$J_{2}$ $=v_{a}\Big{[}iv_{a}+1-x^{2}-2\Delta x^{2}+2x(1-x^{2})\partial_{x}\Big{]}.$ (A1)
In order to obtain Eq. (7) of the main text, we first split $J_{1}$ and $J_{2}$ by whether $x^{2}$ is a factor, writing
$x^{2}\partial_{\tau}\hat{\psi}=x^{2}(L_{1}^{a}\hat{\phi}+L_{2}^{a}\hat{\psi})+J_{1}^{b}\hat{\phi}+J_{2}^{b}\hat{\psi},$ (A2)
in which we define the new operators as
$L_{1}^{a}$ $=v_{a}^{2}\Big{[}\tilde{V}-(1+\Delta)\Delta-2(1+\Delta)x\partial_{x}+(2-x^{2})\partial_{x}^{2}\Big{]},$ $L_{2}^{a}=v_{a}\Big{[}{-}1-2\Delta-2x\partial_{x}\Big{]},$
$J_{1}^{b}$ $=v_{a}^{2}\Big{[}V_{0}+xV_{1}+\Delta+2(1+\Delta)x\partial_{x}-\partial_{x}^{2}\Big{]},$ $J_{2}^{b}=v_{a}\Big{[}iv_{a}+1+2x\partial_{x}\Big{]},\hskip 8.0pt$ (A3)
with $\tilde{V}=(V-V_{0}-xV_{1})/x^{2}$ given by
$\tilde{V}=\sum_{k=0}^{\infty}\frac{x^{k}}{(k+2)!}V_{k+2}.$ (A4)
By writing $\partial_{\tau}\hat{\psi}=\partial_{\tau}\hat{\psi}_{a}+\partial_{\tau}\hat{\psi}_{b}$, we can partition the field so that
$\partial_{\tau}\hat{\psi}_{a}$ $=L_{1}^{a}\hat{\phi}+L_{2}^{a}\hat{\psi},$ (A5)
$x^{2}\partial_{\tau}\hat{\psi}_{b}$ $=J_{1}^{b}\hat{\phi}+J_{2}^{b}\hat{\psi},$ (A6)
satisfying Eq. (A2) by construction. We note that Eq. (A5) is already in the desired form of Eq. (7), whereas Eq. (A6) reproduces the form of Eq. (6). Nevertheless, we can express $\partial_{\tau}\hat{\psi}_{b}$ in the form
$\partial_{\tau}\hat{\psi}_{b}=L_{1}^{b}\hat{\phi}+L_{2}^{b}\hat{\psi},$ (A7)
using a Taylor series expansion about $x=0$. In particular, by repeated differentiation of Eq. (A6), we can obtain all spatial derivatives of $\partial_{\tau}\hat{\psi}_{b}$ at $x=0$. In this way, we obtain the operators $L_{1}^{b}$ and $L_{2}^{b}$ via
$L_{1}^{b}\hat{\phi}=\sum_{k=0}^{\infty}\frac{x^{k}}{(k+2)!}\int^{+1}_{-1}\!\!\text{d}x\ \!\delta(x)(\partial_{x}^{n+2}J_{1}^{b}\hat{\phi}),\quad L_{2}^{b}\hat{\psi}=\sum_{k=0}^{\infty}\frac{x^{k}}{(k+2)!}\int^{+1}_{-1}\!\!\text{d}x\ \!\delta(x)(\partial_{x}^{n+2}J_{2}^{b}\hat{\psi}),$ (A8)
where integration with the Dirac delta function evaluates the derivatives at $x=0$. Explicitly, we find
[…]
$L_{1}=L_{1}^{a}+L_{1}^{b},\quad L_{2}=L_{2}^{a}+L_{2}^{b},$ (A10)
which provides the operators appearing in Eq. (7) of the main text.
A.2 2. Determinant of the Matrix $L^{N}-\Omega\text{Id}$
The determinant of the matrix $L^{N}-\Omega\text{Id}$ may be written as
$\det(L^{N}-\Omega\text{Id})=\det\begin{pmatrix}-\Omega\text{Id}&i\text{Id}\\ iL_{1}^{N}&iL_{2}^{N}-\Omega\text{Id}\end{pmatrix}=\det(\Omega^{2}\text{Id}-i\Omega L_{2}^{N}-L_{1}^{N}),$ (A11)
where we have used that $\Omega\text{Id}$ is invertible.
The QNM spectrum is given when this determinant vanishes, meaning we can scale $\Omega^{2}\text{Id}-i\Omega L_{2}^{N}-L_{1}^{N}$ without changing our result. First, we substitute for the discretized operators $L_{1}^{N}$ and $L_{2}^{N}$, making use of the asymptotic dispersion relation to rewrite $v_{a}$ and $v_{p}$ in terms of the real and imaginary parts of $\Omega$. Then, after an overall scaling of the matrix, we arrive at the condition $\det(M)=0$, with
$M$ $=\tilde{V}^{N}-\sqrt{\Omega}(\sqrt{\Omega}+1)I^{N}+2(\sqrt{\Omega}+1)X^{N}D^{N}+\big{(}2-(X^{N})^{2}\big{)}(D^{N})^{2}$
$\quad\quad+\sum_{k=0}^{N-2}\frac{(X^{N})^{k}}{(k+2)!}\delta^{N}\Big{[}(k+2)V_{1}(D^{N})^{k+1}+[V_{0}+(\sqrt{\Omega}+2(k+2))(\sqrt{\Omega}+1)](D^{N})^{k+2}-(D^{N})^{k+4}\Big{]},$ (A12)
where $X^{N}$ is a diagonal matrix of Chebyshev extremal points and $D^{N}$ is the corresponding Chebyshev differentiation matrix. $\tilde{V}^{N}$ is a diagonal matrix of $\tilde{V}$ evaluated at those same points, and $\delta^{N}$ is the unique matrix representation of integration with the Dirac delta function, obtained by considering its effect on polynomial interpolants.
We see from the form of Eq. (A12) that $\det(M)$ is a polynomial of degree $2N$ in $\sqrt{\Omega}$, with real coefficients, meaning that the solutions $\sqrt{\Omega}$ come in complex conjugate pairs.